### v2

# Exactly solvable model of the square-root price impact dynamics under the long-range market-order correlation

Abstract
In econophysics, there are several enigmatic empirical laws: (i) the market-order flow has strong persistence (long-range order-sign correlation), well formulated as the Lillo-Mike-Farmer model. This phenomenon seems paradoxical given the diffusive and unpredictable price dynamics; (ii) the price impact $I(Q)$ of a large metaorder $Q$ follows the square-root law, $I(Q)\propto\sqrt{Q}$. In this Letter, we propose an exactly solvable model of the nonlinear price-impact dynamics that unifies these enigmas. We generalize the Lillo-Mike-Farmer model to nonlinear price-impact dynamics, which is mapped to an exactly solvable Lévy-walk model. Our exact solution and numerical simulations reveal three important points: First, the price dynamics remains diffusive under the square-root law, even under the long-range correlation. Second, price-movement statistics follows truncated power laws with typical exponent around three. Third, volatility has long memory. While this simple model lacks adjustable free parameters, it naturally aligns even with other enigmatic empirical laws, such as (iii) the inverse-cubic law for price statistics and (iv) volatility clustering. This work illustrates the crucial role of the square-root law in understanding rich and complex financial price dynamics from a single coherent viewpoint.
econophysics, power-law scalings, the square-root law, the Lillo-Mike-Farmer model, Lévy walks
Introduction.
In econophysics, there are four empirical enigmas [1, 2, 3]. The first enigma is the long-range correlation (LRC) of the market-order flow. Market order is an immediate decision to buy or sell stocks at the best prices. By representing the buy (sell) market order at time $t$ by $\epsilon_{t}=+1$ ($\epsilon_{t}=-1$), the LRC states that the autocorrelation (ACF) of the market-order signs exhibits a long memory
$\langle\epsilon_{t}\epsilon_{t+\tau}\rangle_{\rm ss}\propto\tau^{-\gamma},\>\>\>\gamma\in(0,1),$ (1)
where $\langle\dots\rangle_{\rm ss}$ represents the ensemble average in the steady state. This phenomenon originates from the metaorder splitters and is well formulated by the Lillo-Mike-Farmer (LMF) model [5, 4]. Particularly, LMF assumed the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ for the metaorder size $Q$ and predicted the formula
$\gamma=\alpha-1,\>\>\>\alpha\in(1,2),$ (2)
connecting the financial market microstructure and the macroscopic statistics. Recently, the authors validated this prediction quantitatively by scrutinizing the microscopic dataset on the Tokyo Stock Exchange (TSE) [6, 7]. However, this phenomenon is very counter-intuitive: given that the market-order flow is predictable, why is the price dynamics unpredictable? This question is evident given that the price dynamics should exhibit predictable anomalous diffusion if the price impact is linear regarding the metaorder, as documented in the linear propagator model [1, 8, 9] (except under the fine-tuned balance condition between the LRC and the memory). Why is the price dynamics diffusive in the presence of the LRC? This is the first empirical enigma.
The second enigma is the nonlinearity of the price impact, called the square-root law (SRL) [1, 10]. In practice, it has been reported that the price impact is a nonlinear function of the metaorder size $Q$, such that $I(Q):=\langle\epsilon\Delta m\rangle\propto Q^{\delta}$ with $\delta\approx 1/2$, where $\Delta m$ is the price impact by the metaorder $Q$. Recently, the authors scrutinized the TSE dataset, provided a very accurate estimation of $\delta$, and established the strict universality of the SRL, such that $\delta$ exactly equals to $1/2$ for all liquid stocks on the TSE within statistical errors [11]. Still now, the cause of the strict universality of the SRL is unclear. One of the most promising models might be the nonlinear propagator model (called the latent-order book (LLOB) model [1, 10, 12]), but a partially negative evidence was observed on the LLOB hypothesis [13]. Thus, there is no consensus yet regarding the microscopic origin of the SRL.
The third and fourth enigmas are the inverse-cubic law for the price statistics [1, 3, 14, 15, 16, 17, 18] and the volatility clustering [3, 2]: the price movement obeys a fat-tail distribution and the volatility has the long memory:
$P(\Delta m)\propto(\Delta m)^{-\beta-1},\>\>\>\beta\approx 3,$ (3)
$C_{V}(\tau):=$ $\frac{\langle\sigma^{2}(0,t)\sigma^{2}(\tau,t+\tau)\rangle_{\rm ss}}{\langle\sigma^{4}(0,t)\rangle_{\rm ss}}\propto\tau^{-\zeta},\>\>\>\zeta\approx 0.5,$ (4)
where $\sigma^{2}(t,t+\tau):=\{p(t+\tau)-p(t)\}^{2}$. These power laws are ubiquitously observed worldwide independently of asset classes and have been considered as an important part of stylized facts [2, 3]. Despite their long research history, the microscopic origin of these enigmatic laws has been unclear.
Figure 1: Model schematic. Assuming $M$ order splitters with a metaorder size distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ and a rest time distribution $\psi_{r}(\Delta t)=e^{-\Delta t/\tau_{r}}/\tau_{r}$, the price-impact contribution $\Delta m^{(i)}$ for trader $i$ follows the nonlinear scaling $I(Q)\propto Q^{\delta}$. The total price movement $\Delta m$ results from the independent accumulation of all traders’ contributions.
In this Letter, we propose a minimal model that naturally extends the LMF framework to incorporate nonlinear price-impact dynamics, thereby unifying these four enigmas from a single, coherent perspective. We consider the system composed of $M$ order splitters whose the metaorder-size statistics obeys the power law $P(Q)\propto Q^{-\alpha-1}$ with $\alpha\in(1,2)$. All traders are assumed to execute their metaorders whose price impact obeys the nonlinear scaling $I(Q)\propto Q^{\delta}$ with $\delta\in(0,1)$. Crucially, this model is exactly solvable. The price-impact contribution by a single trader can be mapped to the Lévy-walk theory [19, 20] with nonlinear walking-speed down [21]. Our exact solution and numerical simulation have three surprising implications: First, the price dynamics is always diffusive even in the presence the LRC if $\delta=1/2$. Second, the price-movement statistics obeys the power law with $\beta=\alpha/\delta$. This result is consistent with the inverse-cubic law because the typical empirical value of $\alpha$ is $3/2$. Third, we numerically find that our model naturally exhibits the volatility clustering. Thus, our simple model is very simple but minimally consistent with the four important enigmas in econophysics with clear exact solutions.
Model.
The staring point of our model is the LMF model composed of statistically-independent $M$ order splitters with metaorder size $Q$. The size $Q$ is assumed to obey the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $\alpha\in(1,2)$, and the set of the order splitters is denoted by $\Omega_{\rm TR}:=\{1,2,\dots,M\}$ with a positive integer $M>0$. In this Letter, we keep the essentially-identical setup as for the metaorder splitters.
Here we additionally consider the price dynamics triggered by such metaorder splittings (see Fig. 1 for schematic): At $t=0$, trader $i$ waits for the start of metaorder execution according to the exponential resting-time distribution $\psi_{r}(\Delta t)=(1/\tau_{r})e^{-\Delta t/\tau_{r}}$, where $\tau_{r}$ is the average resting time $\tau_{r}$. Then, trader $i$ starts a metaorder execution at the initial time $t^{(i)}_{\rm ini}$ and stops at the final time $t^{(i)}_{\rm fin}$. Here we assume that the price impact exactly obeys the nonlinear price impact $I(Q)\propto Q^{\delta}$ with $\delta\in(0,1)$. For simplicity, we assume that the metaorder time interval $\Delta t^{(i)}:=t_{\rm fin}^{(i)}-t_{\rm ini}^{(i)}$ is proportional to the metaorder size $Q^{(i)}$, such that $Q^{(i)}=\nu\Delta t^{(i)}$, where the executed volume rate $\nu>0$ is an identical constant among traders. The resulting metaorder price impact $\Delta m^{(i)}:=m^{(i)}(t_{\rm fin}^{(i)})-m^{(i)}(t_{\rm ini}^{(i)})$ is assumed to obey the nonlinear scaling: $\Delta m^{(i)}=c\epsilon^{(i)}\left(Q^{(i)}\right)^{\delta}$, where $c>0$ is a constant, $\epsilon^{(i)}$ is the order sign of the trader $i$’s metaorder, and $Q^{(i)}$ is the corresponding metaorder size.
After finalizing the metaorder execution, trader $i$ randomly resets both order sign $\epsilon^{(i)}=\pm 1$ and metaorder size $Q^{(i)}$, takes a rest according to the resting-time distribution $\psi_{r}(\Delta t)$, and then restart his next metaorder execution. For simplicity, we assume $\nu=1$ and $c=1$ by setting the appropriate time and price units.
In this Letter, we consider this deterministic price-impact case as the minimal assumption, since incorporating Gaussian fluctuations has only a minor qualitative effect. Additionally, the volume $Q$ is assumed to be a real number instead of integers for analytical simplicity.
We assume that all traders execute their metaorders statistically independently and that their price-impact contributions independently accumulates. In other words, the price movement $\Delta m(t):=m(t)-m(0)$ is given by
$\Delta m(t)$ $:=\sum_{i\in\Omega_{\rm TR}}\sum_{k=1}^{N^{(i)}(t)}\epsilon_{k}^{(i)}\left(Q_{k}^{(i)}(t)\right)^{\delta},$ (5a)
$Q_{k}^{(i)}(t)$ $:=\min\left\{t_{k;\rm fin}^{(i)},t\right\}-t_{k;\rm ini}^{(i)},$ (5b)
where $t_{k;\rm ini}^{(i)}$ and $t_{k;\rm fin}^{(i)}$ are the initial and final times of the $k$-th metaorder of trader $i$, and $N^{(i)}(t)$ is the total number of metaorders during $[0,t)$ for trader $i$. Also, the order sign $\epsilon_{k}^{(i)}=\pm 1$ is randomly selected with equal probability, and the final metaorder size $Q_{k}^{(i)}=t_{k;\rm fin}^{(i)}-t_{k;\rm ini}^{(i)}$ obeys the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$. The resting time $\Delta t_{k;\rm rest}:=t_{k+1;\rm ini}^{(i)}-t_{k;\rm fin}^{(i)}$ obeys the exponential resting-time distribution $\psi_{r}(\Delta t)=(1/\tau_{r})e^{-\Delta t/\tau_{r}}$. This is the minimal extension of the LMF model by incorporating the nonlinear price impact, by allowing simultaneous metaorder executions among traders on the continuous time axis.
Exact solution.
This model is exactly solvable because the price-impact contribution by a single trader can be mapped to the Lévy-walk theory [19, 20] with nonlinear walking speed [21] (see Appendix A). Indeed, let us define the price-impact contribution by trader $i$ as
$\Delta m^{(i)}(t):=\sum_{k=1}^{N^{(i)}(t)}\epsilon_{k}^{(i)}\left(\Delta t_{k}^{(i)}(t)\right)^{\delta}$ (6)
with $\Delta t_{k}^{(i)}:=\min\{t_{k;\rm fin}^{(i)},t\}-t_{k;\rm ini}^{(i)}$. By interpreting $\Delta t_{k}^{(i)}$ as the “flight time” in Lévy walks, $\Delta m^{(i)}(t)$ represents the cumulative displacement of a Lévy-walk particle with rests and nonlinear space-time coupling. Due to independent price-impact accumulation, the model is solved for any $M>0$.
Diffusive price dynamics under the square-root law.
Figure 2: (a) Phase diagram between superdiffusion and normal diffusion. The phase boundary is given by $2\delta=\alpha$, implying that the price dynamics is always diffusive by assuming the SRL $\delta=1/2$. (b) Numerical mean-squared displacement for $\delta\in\{0.25,0.5\}$ and $\alpha\in\{1.25,1.5,1.75\}$, showing a consistent behavior with our phase diagram. (c) Numerical mean-squared displacement for $\delta\in\{0.75,1\}$, showing the crossover between superdiffusion (orange markers) and normal diffusion (blue markers). (d) Complementary cumulative distribution function for the price changes, showing the inverse-cubic law $P_{>}(\Delta m):=\int_{\Delta m}^{\infty}P(|\Delta m^{\prime}|){\rm d}\Delta m^{\prime}\propto(\Delta m)^{-\beta}$ with $\beta=3$. We assumed the following parameters: $\delta=1/2$, $\alpha=1.5$, $\tau_{r}=10^{5}$, $t=10^{4}$, and $M=1$. See the inset for the corresponding semi-log plot.
Let us present our first main result. The exact solution for the mean-squared displacement is given by
$\langle\Delta m^{2}(t)\rangle\propto\begin{cases}t^{1+2\delta-\alpha}&\mbox{if$2\delta>\alpha$}\\ t&\mbox{if $2\delta<\alpha$}\end{cases},$ (7)
implying that superdiffusion arises if and only if $2\delta>\alpha$. Under the standard LMF assumption $\alpha\in(1,2)$, the price dynamics always exhibits normal diffusion for $\delta\leq 1/2$ (see Fig. 2 (a-c)). See Appendix B for the details.
This result is surprising because it guarantees price diffusion even in the presence of the LRC, assuming the SRL $\delta=1/2$. In other words, thanks to the sufficiently concave nature of the SRL, the market exhibits strong resilience against large liquidity consumption by order splitters. This interpretation is crucially important: it is the SRL that suppresses the large price movement in the presence of the LRC. This scenario highlights the importance of studying the microscopic origin of the SRL for stable regulation of financial markets.
Inverse-cubic law.
We next study the PDF of the price impact $\Delta m$. By evaluating the asymptotic tail of the PDF, we obtain
$P(\Delta m)\propto(\Delta m)^{-\beta-1},\>\>\>\beta:=\frac{\alpha}{\delta}$ (8)
up to the cutoff. This result is clearly consistent with the inverse-cubic law for $\delta=1/2$ because the typical value of $\alpha$ is $3/2$ according to Ref. [5, 22, 6, 7] (see Fig. 2(d)).
For $\beta\leq 2$, the tail can be widely observed for large $t$ because the generalized central limit theorem applies (see Appendix C for its derivation based on the large-deviation principle [23]). On the other hand, for $\beta>2$, the power-law tail regime might not be easily observed depending on the model parameters, because the power-law tail is expected to gradually shrink for large $t$ due to the conventional central limit theorem. However, we identify the parameter regime whereby the power-law tail is clearly observed even for $\beta>2$: When the average resting time satisfies $\tau^{*}\gtrsim t$ with the observation time $t$, we can clearly observe the power-law regime. See Appendix D for the asymptotic evaluation based on the assumption $\tau^{*}\gtrsim t$ (see also Appendix E for a simplified heuristic derivation). This parameter regime is realistic; conversely, it is unrealistic to assume that a splitter rapidly buys and sells huge volumes of stocks within five minutes, given that the inverse cubic law was reported for $t\approx 5$ mins. in [16].
Previously, several theoretical models have been proposed to explain the inverse-cubic law. The first explanation was based on a traditional economic theory [14, 15], one of the main predictions of which was rejected in our previous Letter recently [11]. Some researchers argue that price dynamics follow the Kesten process due to trend-following behavior (see the dealer model [24, 25, 26]), while others suggest that the self-exciting nature of order flows plays a crucial role (see the nonlinear Hawkes process [27, 28, 29], essentially belonging to a nonlinear Kesten family). However, a clear justification has been missing for why $\beta\approx 3$ typically. While our model does not incorporate realistic intraday patterns or external events—both of which we consider essential for accurately predicting $\beta$—it offers a simple and plausible mechanism for why $\beta\approx 3$ is commonly observed.
Volatility clustering.
Figure 3: Volatility ACF in our model, numerically showing the long memory $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$. The model parameter was given by $M=1$, $\delta=0.5$, $\tau_{r}=10$, and $t=1000$. Empirically, the power-law exponent obeys $\zeta\approx\alpha-1$.
Furthermore, we numerically find that this model exhibits the volatility clustering (see Fig. 3): i.e., the ACF of the volatility has the long memory, such that
$C_{V}(\tau)\propto\tau^{-\zeta},\>\>\>\zeta\in(0,1),$ (9)
by assuming the SRL $\delta=1/2$ and $\alpha\in(1,2)$. While we have no mathematical derivation yet, we numerically conjecture an empirical relation
$\zeta\approx\alpha-1.$ (10)
We leave its mathematical derivation for future works.
Concluding discussion.
We develop an exactly solvable model of nonlinear price-impact dynamics in the presence of metaorder splitters. By mapping this model to continuous-time random walks, we analytically derive the statistics of the price dynamics. Assuming the square-root law, our exact solutions demonstrate that the price dynamics remains diffusive, even though the order flow is easily predictable. Furthermore, this model naturally exhibits both inverse-cubic law and volatility clustering. These results are surprising, given that our model, which relies solely on the plausible assumptions of metaorder splitting and the SRL, lacks any trivial mechanisms for replicating both inverse-cubic law and volatility clustering. It is a minimal model, free of artificial memory functions or time-dependent external parameters. While our model is very simple, it provides a unified view on the complex financial price dynamics through the exact solution. It would be interesting to scrutinize the exact solutions by extending our model toward more realistic setups, by introducing (i) stochastic price-impact contributions during metaorder splitting and (ii) the heterogeneity of agents.
Here we discuss the implication of our theory. First, our model is available for various numerical statistical estimation regarding the square-root law. Actually, our model was essentially inspired by the numerical statistical model introduced in our previous Letter [11]. In Ref. [11], we estimated the statistical errors of the estimated $\delta$ by studying a numerical price-impact model exactly obeying the SRL: (i) The price dynamics is essentially identical to the rule (5) by adding stochastic contributions. Other conditions are based on the TSE dataset, such that (ii) the metaorder size $Q_{k}^{(i)}$, the starting time $t_{k;\rm ini}^{(i)}$, and the final time $t_{k;\rm fin}^{(i)}$ of the metaorder executions are identical to the TSE dataset, and (ii) the order signs $\epsilon_{k}^{(i)}$ are randomly shuffled to repeat Monte Carlo simulations. We used such a model to numerically study the consistency and unbiasedness of the statistical estimators therein. While we noticed that the the price dynamics in the numerical model was diffusive at that time, its clear reason was elusive. This Letter provides the theoretical reason why our numerical statistical model provided a plausible time-series even in the presence of the metaorder splitters.
Second, our theory clarifies the practical importance of the SRL regarding the market stability. The presence of the LRC implies that markets suffer from the large liquidity consumption by order splitters. It has been mysterious why markets operate consistently with the unpredictable diffusive nature even in the presence of such a large demand. Our theory shows that the concavity of the SRL alleviates the large liquidity consumption by order splitters to lead the diffusive price dynamics. In other words, if the price impact were less concave than the SRL, the price should have been superdiffusive with clear predictability. As a next step, it would be an interesting to study the statistical characters of the coefficient $c$ in the SRL $I(Q)\propto c\sqrt{Q}$ because it characterises the market resilience against the large demand by large investors.
Acknowledgements.
We thank for the fruitful comments by J.-P. Bouchaud and M. Guillaume. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). KK was supported by JSPS KAKENHI (Grant Nos. 21H01560 and 22H01141) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001). YS developed the numerical program code and contributed to the analytical calculations for the mean-squared displacement. KK designed and managed the project, and contributed the analytical calculations, particularly for the inverse-cubic law. Both YS and KK contributed to writing the manuscript and approved its final content.
Appendix A Appendix
In this Appendix, the outline of our theoretical calculations is given. Here we simplify the notation based on Ref. [19], by rewriting a single-trader contribution $\Delta m^{(j)}(t)$ as the displacement of a Levy-walk particle $x(t)$.
A.1 Appendix A: the Lévy-walk theory
Let us consider a single particle obeying Lévy walks with the flight-time PDF $\psi_{m}(t)=\alpha t^{-\alpha-1}\Theta(t-1)$, the corresponding flight distance $\Delta x=t^{\delta}$, and the exponential resting-time distribution $\psi_{r}(t)=(1/\tau_{r})e^{-t/\tau_{r}}$. The space-time coupling function is given by $\psi(x,t):=(1/2)\delta(|x|-t^{\delta})\psi_{m}(t)$. The PDF of the particle displacement $x$ at time $t$ is written as $P(x,t)$, and its Fourier-Laplace representation is written as $P(k,s):=\int_{0}^{\infty}{\rm d}te^{-ts}\int_{-\infty}^{\infty}{\rm d}xe^{-ts-{\rm i}kx}P(x,t)$. In the following, the arguments $k$ and $s$ always signify the Fourier-Laplace representation.
Assume that $\eta(x,t)$ is the PDF that the particle completes a flight (or equivalently a metaorder splitting) just at time $t$ and arrives at $x$. This conditional PDF $\eta(x,t)$ satisfies a recursive relation
$\eta(x,t)=\delta(x)\delta(t)+$ (11)
$\int{\rm d}x_{1}\int_{0}^{t}{\rm d}t_{1}\int_{t_{1}}^{t}{\rm d}t_{2}\eta(x_{1},t_{1})\psi_{r}(t_{2}-t_{1})\psi(x-x_{1},t-t_{2}).$
Using $\eta(x,t)$, the PDF $P(x,t)$—the probability density function that the particle resides at the position $x$ at time $t$ (without conditioning of the flight completion)—is given by
$P(x,t)=\int{\rm d}x_{1}\int_{0}^{t}{\rm d}t_{1}\eta(x_{1},t_{1})\Psi_{r}(t-t_{1})+$ (12)
$\int{\rm d}x_{1}\int_{0}^{t}{\rm d}t_{1}\int_{t_{1}}^{t}{\rm d}t_{2}\eta(x_{1},t_{1})\psi_{r}(t_{2}-t_{1})\Psi(x-x_{1},t-t_{2}),$
where $\Psi_{r}(t):=\int_{t}^{\infty}\psi_{r}(t^{\prime}){\rm d}t^{\prime}$, $\Psi_{m}(t):=\int_{t}^{\infty}\psi_{m}(t^{\prime}){\rm d}t^{\prime}$, and $\Psi(x,t):=(1/2)\delta(|x|-t^{\delta})\Psi_{m}(t)$. By applying the Fourier-Laplace transform, we obtain
$P(k,s)$ $=\frac{\Psi_{r}(s)+\Psi(k,s)\psi_{r}(s)}{1-\psi_{r}(s)\psi(k,s)},$ (13)
$\Psi(k,s)$ $=\int^{\infty}_{0}{\rm d}t\Psi_{m}(t)e^{-st}\cos(kt^{\delta}),$ (14)
where $\psi_{r}(s)=1/(1+\tau_{r}s)$ and $\Psi_{r}(s)=\tau_{r}/(1+\tau_{r}s)$. This is the exact PDF for the contribution by a single trader.
While we focused on the market-impact contribution by a single trader, it is straightforward to study the market-impact contribution by multiple traders. Indeed, by writing the total market impact as $x_{\rm tot}:=\sum_{i=1}^{M}x^{(i)}$ with the $i$-th trader’s contribution $x^{(i)}$, we obtain its characteristic function as
$\left\langle e^{-{\rm i}kx_{\rm tot}}\right\rangle=\prod_{i=1}^{M}\left\langle e^{-{\rm i}kx^{(i)}}\right\rangle=\left[\left\langle e^{-{\rm i}kx^{(i)}}\right\rangle\right]^{M}$ (15)
because $x^{(i)}$ is independent of $x^{(j)}$ for $i\neq j$. Thus, studying a single-trader market impact is sufficient to understand the statistics of the total market impact. For example, we obtain $\langle x^{2}_{\rm tot}\rangle=M\langle(x^{(i)})^{2}\rangle$ because $\langle x^{(i)}\rangle=0$.
A.2 Appendix B: the mean-squared displacement
Let us expand $P(k,s)$ regarding $k$ by fixing $s$. From Eq. (13), we obtain
$P(k,s)\approx s^{-1}-\frac{k^{2}}{2}\left(As^{\alpha-2\delta-2}+Bs^{-2}+o(s^{-2})\right)+o(k^{2})$ (16)
with some coefficients $A$ and $B$. By applying the inverse Laplace transform, we obtain the asymptotic behaviour (7) of the mean-squared displacement.
A.3 Appendix C: the inverse-cubic law based on the large-deviation principle
Let us derive the power-law statistics for the price movement by assuming the large-deviation principle [23] for large $t$:
$P(k,t):=\int_{0}^{\infty}P(x,t)e^{-{\rm i}kx}{\rm d}x\approx e^{-t\Lambda(k)+o(t)},$ (17)
where $\Lambda(s)$ is the cumulant generating function function. We can evaluate its Laplace representation $P(k,s)$ as $P(k,s)=\int_{0}^{\infty}e^{-st-t\Lambda(k)}{\rm d}t\approx 1/[s+\Lambda(k)]$ for small $s$. We thus obtain the cumulant generating function by the formula
$\Lambda(k)=\lim_{s\to 0}\frac{1}{P(k,s)}=\frac{1-\psi(k,s=0)}{\tau_{r}+\Psi(k,s=0)}$ (18)
By expanding $\psi(k,s=0)$ and $\Psi(k,s=0)$ for small $k$, we obtain
$\Lambda(k)\approx-A^{\prime}k^{2}-B^{\prime}|k|^{\beta}+o(|k|^{\beta})+o(s),\>\>\>\beta:=\frac{\alpha}{\delta}$ (19)
with some coefficients $A^{\prime}$ and $B^{\prime}$. By using the Tauberian theorem [19, 29], we obtain the power-law tail (8).
A.4 Appendix D: the inverse-cubic law based on a long resting-time approximation
Let us consider the case where the average resting time is sufficiently large $\tau_{r}\gg t$. Under this assumption, we can expand the exact solution (13) to obtain
$P(k,s)\approx$ $\Psi_{r}(s)+\Psi(k,s)\psi_{r}(s)$
$+\Psi_{r}(s)\psi(k,s)\psi_{r}(s)+O(\tau_{r}^{-2}),$ (20)
which leads to a truncated power-law asymptotics
$P(x,t)\approx\frac{t}{2\tau_{r}\delta}|x|^{-1-\beta}\mbox{ for $1\ll t\ll\tau_{r}$}$ (21)
by focusing on the regime $1\ll|x|\ll t^{\delta}$.
A.5 Appendix E: the inverse-cubic law based on heuristic arguments
Let us roughly assume that the price movement $x$ is proportional to the largest flight-jump size $t^{\delta}$. For the power-law PDF $\psi_{m}(t)\propto t^{-1-\alpha}$, by the Jacobian relation
$\psi_{m}(t){\rm d}t=P(x){\rm d}x$ (22)
with $x\propto t^{\delta}$, we obtain
$P(x)\propto x^{-\beta-1},\>\>\>\beta:=\frac{\alpha}{\delta}.$ (23)
This simplified derivation essentially captures the picture under the long resting-time approximation, where the total displacement is primarily determined by a single largest jump.