### v4

# Why do financial prices exhibit Brownian motion despite predictable order flow?

Abstract
In econophysics of financial market microstructure, there are two enigmatic empirical laws: (i) the market-order flow has predictable persistence due to metaorder splitters by institutional investors, well formulated as the Lillo-Mike-Farmer model. However, this phenomenon seems paradoxical given the diffusive and unpredictable price dynamics; (ii) the price impact $I(Q)$ of a large metaorder $Q$ follows the square-root law, $I(Q)\propto\sqrt{Q}$. Here we theoretically reveal why price dynamics follows Brownian motion despite predictable order flow by unifying these enigmas. We generalise the Lillo-Mike-Farmer model to nonlinear price-impact dynamics, which is mapped to an exactly solvable Lévy-walk model. Our exact solution shows that the price dynamics remains diffusive under the square-root law, even under persistent order flow. This work illustrates the crucial role of the square-root law in mitigating large price movements by large metaorders, thereby leading to the Brownian price dynamics, consistent with the efficient market hypothesis over long timescales.
econophysics, market microstructure, the square-root law, the Lillo-Mike-Farmer model, Lévy walks
Figure 1: Order-splitting behaviour (left) and the corresponding price impact (right). Institutional traders typically split their large metaorders into a long sequences of small child orders to minimize their transaction cost. The long-term predictability (long memory) of market-order flow arises due to this order splitting because the order signs of the child orders are identical. The resulting market impact—the price change between the initial and final times of metaorders—empirically follows a nonlinear scaling $I(Q)\propto\sqrt{Q}$, called the square-root law.
Introduction.
There are two empirical enigmas in econophysics [1, 2, 3]. The first enigma is the long memory of the market-order flow. A market order is an immediate decision to buy or sell stocks, which is represented by $\epsilon_{t}=+1$ ($\epsilon_{t}=-1$) for the buy (sell) market order at time $t$ (e.g., $\{\epsilon_{t}\}_{t\geq 1}=\{+,+,+,-,+\}$ signifies three buys, one sell, and one buy orders). The long memory implies that such order signs are easily predictable and, thus, the autocorrelation (ACF) of the market-order signs decays very slowly:
[…]
where $E[\dots]$ represents the ensemble average in the steady state and the ACF’s sum diverges as $\sum_{\tau=1}^{\infty}E[\epsilon_{t}\epsilon_{t+\tau}]=\infty$. This formula implies that the order flow has persistence for a very long time, such as for a few days [1].
Then, why does this predictability arise ubiquitously in financial markets? This order-sign predictability originates from the metaorder splitters and is well formulated by the Lillo-Mike-Farmer (LMF) model [4, 5]. The LMF theory argues that institutional investors have hidden large orders (called metaorders $Q$), which are split into a series of child orders. Since the order signs of the child orders should be identical, the order flow has natural predictability (see Fig. 1). Particularly, the LMF theory predicted a formula connecting the market microstructure and the ACF:
[…]
where the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ is assumed for the metaorder size $Q$. Recently, the authors validated this prediction quantitatively by scrutinizing the microscopic dataset of the Tokyo Stock Exchange (TSE) [6, 7] and, thereby, have established the LMF theory as the causality theory for the long memory.
However, this phenomenon is very counter-intuitive: given that the market-order flow is predictable, why is the price dynamics unpredictable? This puzzle arises because the price dynamics would exhibit predictable anomalous diffusion if the price impact were linear regarding the metaorder size, as documented in the linear propagator model [1, 8, 9] (except under the fine-tuned balance condition between the order-flow and price-impact decays). Why is the price dynamics diffusive despite the predictable order flow? This is the first empirical enigma, particularly because most traditional economic theory predicts linear price-impact models [10].
The second enigma is the nonlinearity of the price impact, called the square-root law (SRL) [1, 11] (see Fig. 1). In practice, the linear price-impact model is valid only for small metaorder size $Q$; for large $Q$, it has been reported that the price impact is a nonlinear function of the metaorder size $Q$:
$I(Q):=E[\epsilon\Delta m\mid Q]\propto Q^{\delta},\>\>\>\delta\approx\frac{1}{2},$ (3)
where $\Delta m$ is the price impact by the metaorder $Q$. Recently, the authors scrutinized the TSE dataset, provided a very accurate estimation of $\delta$, and established the strict universality of the SRL, such that $\delta$ exactly equals to $1/2$ for all liquid stocks on the TSE within statistical errors [12]. Still now, the cause of the strict universality of the SRL is unclear. One of the most promising models might be the nonlinear propagator model (called the latent-order book model [1, 11, 13]), but a partially negative evidence was observed on the latent-order book model [14]. Thus, there is no consensus yet regarding the microscopic origin of the SRL.
Figure 2: Model schematic. Assuming $M$ order splitters with a metaorder size distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ and a rest time distribution $\psi_{r}(\Delta t)=e^{-\Delta t/\tau_{r}}/\tau_{r}$, the price-impact contribution $\Delta m^{(i)}$ for trader $i$ follows the nonlinear scaling $I(Q)\propto Q^{\delta}$. The total price movement $\Delta m$ results from the independent accumulation of all traders’ contributions.
Here we propose a minimal model that naturally extends the LMF framework to incorporate nonlinear price-impact dynamics, thereby unifying these two enigmas from a single, coherent perspective. We consider the system composed of $M$ order splitters whose metaorder-size statistics follow the power law $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $21<\alpha<$ 2. All traders are assumed to execute their metaorders whose price impact obeys the nonlinear scaling $I(Q)\propto Q^{\delta}$ with $0<\delta\leq 1$. Crucially, this model is exactly solvable. The price-impact contribution by a single trader can be mapped to the Lévy-walk theory [15, 16] with nonlinear walking-speed down [17]. Our exact solution shows that the price dynamics is always diffusive even in the presence of the long memory if $\delta=1/2$. This result implies that the SRL plays the crucial role in mitigating the price impact due to the huge liquidity consumption by institutional investors. Additionally, we find that this model exhibits consistent behavior with other empirical laws (such as the inverse-cubic law [1, 3, 18, 19, 20, 21, 22] and volatility clustering [3, 2]). Thus, our simple model is very simple but minimally consistent with various empirical enigmas in econophysics of market microstructure with clear exact solutions.
Model.
The staring point of our model is the LMF model composed of statistically-independent $M$ order splitters with metaorder size $Q$. The size $Q$ is assumed to follow the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $21<\alpha<$ 2, and the set of the order splitters is denoted by $\Omega_{\rm TR}:=\{1,2,\dots,M\}$ with a positive integer $M>0$. In this Letter, we keep the essentially-identical setup as for the metaorder splitters.
Here we additionally consider the price dynamics triggered by such metaorder splittings (see Fig. 2 for schematic): At $t=0$, trader $i$ waits for the start of metaorder execution according to the exponential resting-time distribution $\psi_{r}(\Delta t)=(1/\tau_{r})e^{-\Delta t/\tau_{r}}$, where $\tau_{r}$ is the average resting time $\tau_{r}$. Then, trader $i$ starts a metaorder execution at the initial time $t^{(i)}_{\rm ini}$ and stops at the final time $t^{(i)}_{\rm fin}$. Here we assume that the price impact exactly obeys the nonlinear price impact $I(Q)\propto Q^{\delta}$ with $\delta\in(0,1)$. For simplicity, we assume that the metaorder time interval $\Delta t^{(i)}:=t_{\rm fin}^{(i)}-t_{\rm ini}^{(i)}$ is proportional to the metaorder size $Q^{(i)}$, such that $Q^{(i)}=\nu\Delta t^{(i)}$, where the executed volume rate $\nu>0$ is an identical constant among traders. The resulting metaorder price impact $\Delta m^{(i)}:=m^{(i)}(t_{\rm fin}^{(i)})-m^{(i)}(t_{\rm ini}^{(i)})$ is assumed to obey the nonlinear scaling: $\Delta m^{(i)}=c\epsilon^{(i)}\left(Q^{(i)}\right)^{\delta}$, where $c>0$ is a constant, $\epsilon^{(i)}$ is the order sign of the trader $i$’s metaorder, and $Q^{(i)}$ is the corresponding metaorder size.
Upon completing a metaorder, trader $i$ resets both order sign $\epsilon^{(i)}=\pm 1$ and metaorder size $Q^{(i)}$, takes a rest according to the resting-time distribution $\psi_{r}(\Delta t)$, and restart his next metaorder execution. For simplicity, we set $\nu=c=1$ by appropriate choice of time and price units.
In this work, we consider this deterministic price-impact case as the minimal assumption, since incorporating Gaussian fluctuations has only a minor qualitative effect. Additionally, the volume $Q$ is assumed to be a real number instead of integers for analytical simplicity. Also, the impact decay after metaorder splitting is not considered in this work.
We assume that all traders execute their metaorders statistically independently and that their price-impact contributions independently accumulates. In other words, the price movement $\Delta m(t):=m(t)-m(0)$ is given by
$\Delta m(t)$ $:=\sum_{i\in\Omega_{\rm TR}}\sum_{k=1}^{N^{(i)}(t)}\epsilon_{k}^{(i)}\left(Q_{k}^{(i)}(t)\right)^{\delta},$ (4a)
$Q_{k}^{(i)}(t)$ $:=\min\left\{t_{k;\rm fin}^{(i)},t\right\}-t_{k;\rm ini}^{(i)},$ (4b)
where $t_{k;\rm ini}^{(i)}$ and $t_{k;\rm fin}^{(i)}$ are the initial and final times of the $k$-th metaorder of trader $i$, and $N^{(i)}(t)$ is the total number of metaorders during $[0,t)$ for trader $i$. Also, the order sign $\epsilon_{k}^{(i)}=\pm 1$ is randomly selected with equal probability, and the final metaorder size $Q_{k}^{(i)}=t_{k;\rm fin}^{(i)}-t_{k;\rm ini}^{(i)}$ follows the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$. The resting time $\Delta t_{k;\rm rest}:=t_{k+1;\rm ini}^{(i)}-t_{k;\rm fin}^{(i)}$ follows the exponential resting-time distribution $\psi_{r}(\Delta t)=(1/\tau_{r})e^{-\Delta t/\tau_{r}}$. This is the minimal extension of the LMF model by incorporating the nonlinear price impact, by allowing simultaneous metaorder executions among traders on the continuous time axis.
Exact solution.
This model is exactly solvable because the price-impact contribution by a single trader can be mapped to the Lévy-walk theory [15, 16] with nonlinear walking speed [17] (see Methods). Indeed, let us define the price-impact contribution by trader $i$ as
$\Delta m^{(i)}(t):=\sum_{k=1}^{N^{(i)}(t)}\epsilon_{k}^{(i)}\left(\Delta t_{k}^{(i)}(t)\right)^{\delta}$ (5)
with $\Delta t_{k}^{(i)}:=\min\{t_{k;\rm fin}^{(i)},t\}-t_{k;\rm ini}^{(i)}$. By interpreting $\Delta t_{k}^{(i)}$ as the “flight time” in Lévy walks, $\Delta m^{(i)}(t)$ represents the cumulative displacement of a Lévy-walk particle with rests and nonlinear space-time coupling. Due to independent price-impact accumulation, the model is solved for any $M>0$.
Brownian price dynamics under the square-root law.
Figure 3: (a) Phase diagram between superdiffusion and normal diffusion. The phase boundary is given by $2\delta=\alpha$, implying that the price dynamics is always diffusive by assuming the SRL $\delta=1/2$. (b) Numerical mean-squared displacement for $\delta\in\{0.25,0.5\}$, $\alpha\in\{1.25,1.5,1.75\}$, and $N=1$, showing a consistent behavior with our phase diagram. (c) Numerical mean-squared displacement for $\delta\in\{0.75,1\}$, showing the crossover between superdiffusion (orange markers) and normal diffusion (blue markers). More detailed numerical results are presented in Supplementary informationfor other parameters.
Let us present our first main result. The exact solution for the mean-squared displacement is given by
$E[\Delta m^{2}(t)]\propto\begin{cases}t^{1+2\delta-\alpha}&\mbox{if $2\delta>\alpha$}\\ t&\mbox{if $2\delta<\alpha$}\end{cases},$ (6)
implying that superdiffusion arises if and only if $2\delta>\alpha$. Under the standard LMF assumption $21<\alpha<$ 2, the price dynamics always exhibits normal diffusion for $\delta\leq 1/2$ (see Fig. 3). See Methods for the details. Note that this result is consistent with the linear propagator model [8, 9] for the specific case with non-zero permanent impact for $\delta=1$.
This result is surprising because it guarantees price diffusion even in the presence of the long memory in the order flow, assuming the SRL $\delta=1/2$. In other words, thanks to the sufficiently concave nature of the SRL, the market exhibits strong resilience against large liquidity consumption by order splitters. This interpretation is crucially important: it is the SRL that suppresses the large price movement in the presence of the long memory. This scenario highlights the importance of studying the microscopic origin of the SRL for stable regulation of financial markets.
Concluding discussion.
We develop an exactly solvable model of nonlinear price-impact dynamics in the presence of metaorder splitters. By mapping this model to continuous-time random walks, we analytically derive the statistics of the price dynamics. Assuming the SRL, our exact solution demonstrates that the price dynamics remains diffusive, even though the order flow is easily predictable. It would be interesting to scrutinize the exact solutions by extending our model toward more realistic setups, by introducing (i) stochastic price-impact contributions during metaorder splitting and (ii) the heterogeneity of agents.
Here we discuss the implication of our theory. First, our theory clarifies the practical importance of the SRL regarding the market stability. The presence of the long memory implies that markets suffer from the large liquidity consumption by order splitters. It has been mysterious why markets operate consistently with the unpredictable diffusive nature even in the presence of such a large demand. Our theory shows that the concavity of the SRL alleviates the large liquidity consumption by order splitters, which leads to diffusive price dynamics. In other words, if the price impact were less concave than the SRL, the price should have been superdiffusive with clear predictability. As a next step, it would be an interesting topic to study the statistical characters of the coefficient $c$ in the SRL $I(Q)\propto c\sqrt{Q}$ because it characterises the market resilience against the large demand by large investors.
Second, our theory highlights that the SRL is not a technical minor topic but should be regarded as a key topic in financial economics because it is directly related to the theoretical foundation of the efficient market hypothesis (EMH)—the hypothesis stating that the market is sufficiently efficient and the price is unpredictable (thus diffusive) from public information. Evidently, the EMH is the most important concept for various financial studies. However, most of traditional economic theories (such as Kyle’s one [10]) predict the linear price-impact law, which apparently contradicts the EMH in the presence of the predictable market-order flow. Our theory clarifies that, even under the predictable market-order flow, the EMH robustly holds at a long timescale thanks to the SRL. In other words, it is the SRL that prohibits simple arbitrage strategies from professional traders even when the market-order flow has the predictable long memory. There is the room to update financial-economic theories by incorporating the SRL to keep the theoretical consistency between the EMH and the predictable market-order flow.
Third, our model is available for various numerical statistical estimation regarding the SRL. Actually, our model was essentially inspired by the numerical statistical model introduced in our previous Letter [12]. In Ref. [12], we estimated the statistical errors of the estimated $\delta$ by studying a numerical price-impact model exactly obeying the SRL: (i) The price dynamics is essentially identical to the rule (4) by adding stochastic contributions. Other conditions are based on the TSE dataset, such that (ii) the metaorder size $Q_{k}^{(i)}$, the starting time $t_{k;\rm ini}^{(i)}$, and the final time $t_{k;\rm fin}^{(i)}$ of the metaorder executions are identical to the TSE dataset, and (ii) the order signs $\epsilon_{k}^{(i)}$ are randomly shuffled to repeat Monte Carlo simulations. We used such a model to numerically study the consistency and unbiasedness of the statistical estimators therein. While we noticed that the the price dynamics in the numerical model was diffusive at that time, its clear reason was elusive. This Letter provides the theoretical reason why our numerical statistical model provided a plausible time-series even in the presence of the metaorder splitters.
Fourth, interestingly, our simple model is consistent with even other enigmatic empirical laws in finance without fine-tuning. For example, the price change $\Delta m$ follows the inverse-cubic law [1, 3, 18, 19, 20, 21, 22]
$P(\Delta m)\propto(\Delta m)^{-\beta-1},\>\>\>\beta:=\frac{\alpha}{\delta},$ (7a)
clearly consistent with the inverse-cubic law $\beta\approx 3$ for $\delta=1/2$ because the typical value of $\alpha$ is $3/2$ according to Ref. [4, 23, 6, 7] (see Methods). Furthermore, our model exhibits even volatility clustering [3, 2]:
$C_{V}(\tau)$ $:=\frac{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})(\sigma^{2}(\tau,\tau+t)-\mu_{\sigma^{2}_{t}})]}{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})^{2}]}$ (7b)
[…]
where $\mu_{\sigma^{2}_{t}}=E[\sigma^{2}(0,t)]$, and $\sigma^{2}(t,t+\tau):=\{p(t+\tau)-p(t)\}^{2}$, by assuming the SRL $\delta=1/2$. While we have no mathematical derivation, we numerically conjecture $\zeta\approx\alpha-1$ (see Methods and Fig. 4). These two characters (7) are surprising, given that our model, which relies solely on the plausible assumptions of metaorder splitting and the SRL, lacks any trivial mechanisms for replicating both inverse-cubic law and volatility clustering. It is a minimal model, free of artificial memory functions or time-dependent external parameters.
Finally, let us address the potential limitation of our theory. Our model incorporates neither market-microstructure dynamics nor time-dependent parameters in our model—be they endogenous (e.g., self-excitation mechanism [24, 25, 26] and the limit-order book dynamics) or exogenous (e.g., intraday seasonality and external shocks)—, which is both a strength and a weakness. A strength is the model’s parameter-free nature, allowing direct testability with microscopic datasets. A weakness is its inability to flexibly capture realistic endogenous/exogenous market microscopic dynamics, and we are unsure yet if our simple model—accounting for only order-splitting and the SRL—fully captures the causal mechanisms behind the inverse-cubic law and the volatility clustering. Several researchers claim that limit-order-book conditions and liquidity crises are key to understanding the large price jump [27, 28], but these points are missing in our model. Incorporating such mechanisms into our model would be an important next step.
I Methods
In this section, the outline of our theoretical calculations is given. Here we simplify the notation based on Ref. [15], by rewriting a single-trader contribution $\Delta m^{(j)}(t)$ as the displacement of a Levy-walk particle $x(t)$. More detailed calculations are presented in Supplementary information.
I.1 Appendix A: the Lévy-walk theory
Let us consider a single particle obeying Lévy walks with the flight-time PDF $\psi_{m}(t)=\alpha t^{-\alpha-1}\Theta(t-1)$, the corresponding flight distance $\Delta x=t^{\delta}$, and the exponential resting-time distribution $\psi_{r}(t)=(1/\tau_{r})e^{-t/\tau_{r}}$. The space-time coupling function is given by $\psi(x,t):=(1/2)\delta(|x|-t^{\delta})\psi_{m}(t)$. The PDF of the particle displacement $x$ at time $t$ is written as $P(x,t)$, and its Fourier-Laplace representation is written as $P(k,s):=\int_{0}^{\infty}{\rm d}t\int_{-\infty}^{\infty}{\rm d}xe^{-ts-{\rm i}kx}P(x,t)$. In the following, the arguments $k$ and $s$ always signify the Fourier-Laplace representation.
Assume that $\eta(x,t)$ is the PDF that the particle completes a flight (or equivalently a metaorder splitting) just at time $t$ and arrives at $x$. This conditional PDF $\eta(x,t)$ satisfies a recursive relation
$\eta(x,t)=\delta(x)\delta(t)+$ (8)
$\int{\rm d}x_{1}\int_{0}^{t}{\rm d}t_{1}\int_{t_{1}}^{t}{\rm d}t_{2}\eta(x_{1},t_{1})\psi_{r}(t_{2}-t_{1})\psi(x-x_{1},t-t_{2}).$
Using $\eta(x,t)$, the PDF $P(x,t)$—the probability density function that the particle resides at the position $x$ at time $t$ (without conditioning of the flight completion)—is given by
$P(x,t)=\int{\rm d}x_{1}\int_{0}^{t}{\rm d}t_{1}\eta(x_{1},t_{1})\Psi_{r}(t-t_{1})+$ (9)
$\int{\rm d}x_{1}\int_{0}^{t}{\rm d}t_{1}\int_{t_{1}}^{t}{\rm d}t_{2}\eta(x_{1},t_{1})\psi_{r}(t_{2}-t_{1})\Psi(x-x_{1},t-t_{2}),$
where $\Psi_{r}(t):=\int_{t}^{\infty}\psi_{r}(t^{\prime}){\rm d}t^{\prime}$, $\Psi_{m}(t):=\int_{t}^{\infty}\psi_{m}(t^{\prime}){\rm d}t^{\prime}$, and $\Psi(x,t):=(1/2)\delta(|x|-t^{\delta})\Psi_{m}(t)$. By applying the Fourier-Laplace transform, we obtain
$P(k,s)$ $=\frac{\Psi_{r}(s)+\Psi(k,s)\psi_{r}(s)}{1-\psi_{r}(s)\psi(k,s)},$ (10)
$\Psi(k,s)$ $=\int^{\infty}_{0}{\rm d}t\Psi_{m}(t)e^{-st}\cos(kt^{\delta}),$ (11)
where $\psi_{r}(s)=1/(1+\tau_{r}s)$ and $\Psi_{r}(s)=\tau_{r}/(1+\tau_{r}s)$. This is the exact PDF for the contribution by a single trader.
While we focused on the market-impact contribution by a single trader, it is straightforward to study the market-impact contribution by multiple traders. Indeed, by writing the total market impact as $x_{\rm tot}:=\sum_{i=1}^{M}x^{(i)}$ with the $i$-th trader’s contribution $x^{(i)}$, we obtain its characteristic function as
$E\left[e^{-{\rm i}kx_{\rm tot}}\right]=\prod_{i=1}^{M}E\left[e^{-{\rm i}kx^{(i)}}\right]=\left\{E\left[e^{-{\rm i}kx^{(i)}}\right]\right\}^{M}$ (12)
because $x^{(i)}$ is independent of $x^{(j)}$ for $i\neq j$. Thus, studying a single-trader market impact is sufficient to understand the statistics of the total market impact. For example, we obtain $E[x^{2}_{\rm tot}]=M\times E[(x^{(i)})^{2}]$ because $E[x^{(i)}]=0$.
I.2 Appendix B: the mean-squared displacement
Let us expand $P(k,s)$ regarding $k$ by fixing $s$. From Eq. (10), we obtain
$P(k,s)\approx s^{-1}-\frac{k^{2}}{2}\left(As^{\alpha-2\delta-2}+Bs^{-2}+o(s^{-2})\right)+o(k^{2})$ (13)
with some coefficients $A$ and $B$. By applying the inverse Laplace transform, we obtain the asymptotic behaviour (6) of the mean-squared displacement.
I.3 Appendix C: the inverse-cubic law
Figure 4: (a) Complementary cumulative distribution function for the price changes, showing the inverse-cubic law $P_{>}(\Delta m)\propto(\Delta m)^{-\beta}$ with $\beta=3$. We assumed the following parameters: $\delta=1/2$, $\alpha=1.5$, $\tau_{r}=10^{5}$, $t=10^{4}$, and $M=1$. See the inset for the corresponding semi-log plot. (b) Volatility ACF in our model, numerically showing the long memory $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$ for the parameters $M=1$, $\delta=0.5$, $\tau_{r}=1$, and $t=10$. Empirically, the power-law exponent obeys $\zeta\approx\alpha-1$. More detailed numerical investigation is presented in the SI.
Our model exhibits the inverse-cubic law (7a), or equivalently,
$P_{>}(\Delta m):=\int_{\Delta m}^{\infty}P(|\Delta m^{\prime}|){\rm d}\Delta m^{\prime}\propto(\Delta m)^{-\beta},\>\>\>\beta:=\frac{\alpha}{\delta},$ (14)
where $P_{>}(\Delta m)$ is the complementary cumulative distribution function for the price changes $\Delta m:=m(t)-m(0)$. See Fig. 4(a) for the numerical simulation. Let us derive this formula by three methods below.
I.3.1 Derivation based on the large deviation principle
Let us derive the power-law statistics for the price movement by assuming the large-deviation principle [29] for large $t$:
$P(k,t):=\int_{0}^{\infty}P(x,t)e^{-{\rm i}kx}{\rm d}x\approx e^{-t\Lambda(k)+o(t)},$ (15)
where $\Lambda(s)$ is the cumulant generating function function. We can evaluate its Laplace representation $P(k,s)$ as $P(k,s)=\int_{0}^{\infty}e^{-st-t\Lambda(k)}{\rm d}t\approx 1/[s+\Lambda(k)]$ for small $s$. We thus obtain the cumulant generating function by the formula
$\Lambda(k)=\lim_{s\to 0}\frac{1}{P(k,s)}=\frac{1-\psi(k,s=0)}{\tau_{r}+\Psi(k,s=0)}$ (16)
By expanding $\psi(k,s=0)$ and $\Psi(k,s=0)$ for small $k$, we obtain
$\Lambda(k)\approx-A^{\prime}k^{2}-B^{\prime}|k|^{\beta}+o(|k|^{\beta})+o(s),\>\>\>\beta:=\frac{\alpha}{\delta}$ (17)
with some coefficients $A^{\prime}$ and $B^{\prime}$. By using the Tauberian theorem [15, 26], we obtain the power-law tail (7a).
I.3.2 Derivation based on a long resting-time approximation
Let us consider the case where the average resting time is sufficiently large $\tau_{r}\gg t$. Under this assumption, we can expand the exact solution (10) to obtain
$P(k,s)\approx$ $\Psi_{r}(s)+\Psi(k,s)\psi_{r}(s)$
$+\Psi_{r}(s)\psi(k,s)\psi_{r}(s)+O(\tau_{r}^{-2}),$ (18)
which leads to a truncated power-law asymptotics
$P(x,t)\approx\frac{t}{2\tau_{r}\delta}|x|^{-1-\beta}\mbox{ for $1\ll t\ll\tau_{r}$}$ (19)
by focusing on the regime $1\ll|x|\ll t^{\delta}$.
I.3.3 Heuristic derivation
Let us roughly assume that the price movement $x$ is proportional to the largest flight-jump size $t^{\delta}$. For the power-law PDF $\psi_{m}(t)\propto t^{-1-\alpha}$, by the Jacobian relation
$\psi_{m}(t){\rm d}t=P(x){\rm d}x$ (20)
with $x\propto t^{\delta}$, we obtain
$P(x)\propto x^{-\beta-1},\>\>\>\beta:=\frac{\alpha}{\delta}.$ (21)
This simplified derivation essentially captures the picture under the long resting-time approximation, where the total displacement is primarily determined by a single largest jump.
II Appendix D: Volatility clustering
We also numerically study the long memory of the volatility in our model. The model parameter was given by $M=1$, $\delta=1/2$, $\tau_{r}=10$, and $t=1000$. In Fig. 4(b), we plot the volatility ACF
$C_{V}:=\frac{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})(\sigma^{2}(\tau,\tau+t)-\mu_{\sigma^{2}_{t}})]}{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})^{2}]},$ (22)
with $\sigma^{2}(t,t+\tau):=\{p(t+\tau)-p(t)\}^{2}$ and $\mu_{\sigma^{2}_{t}}:=E[\sigma^{2}(0,t)]$. The figure shows the power-law decay $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$. We numerically find the relationship
$\zeta\approx\alpha-1.$ (23)
We leave the mathematical derivation of this empirical formula as a future issue.
Code availability
Simulation codes are available as our SI.
Acknowledgements
Acknowledgements.
We thank for the fruitful comments by J.-P. Bouchaud and M. Guillaume. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). KK was supported by JSPS KAKENHI (Grant Nos. 21H01560 and 22H01141) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001).
Author contribution
YS developed the numerical program code and contributed to the analytical calculations for the mean-squared displacement. KK designed and managed the project, and contributed the analytical calculations, particularly for the inverse-cubic law. Both YS and KK contributed to writing the manuscript and approved its final content.
Conflict of interest
We declare no financial conflict of interest.